ROSAT S/C Status 1998-12-04 ... 1999-01-14

Compiled by Jakob Englhauser
.. HOME .. Up ..
Fri 1998-12-04 (DoY 338)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    HRI is switched off, and in off position.
    WFC is in safemode.

  10:39:00   WHM1  AMCS: steady
                   PSPC: filling from tank B
                      tank / PSPC pressure (mB)
                      ?920 /  50
  12:18:00   WHM2     ?900 / 220
  13:58:00   WHM3      880 / 480
  15:37:30   WHM4      865 / 622


Mon 1998-12-07 (DoY 341)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC is under pressure, and in off position.
    WFC is in safemode.

  08:35:00   WHM1  PSPC: switching from tank B to tank A+C
                      pressure (mB) tank   A /  B  /  C  / PSPC
                                        1857  779  1794     602
                     [current exceeded; valves probably not activated]

  10:12:15   WHM2  PSPC: valve problem fixed; 632 mB

  11:52:00   WHM3  PSPC: 1019 mB; AMCS: steady

  13:32:00   WHM4  PSPC: 
                     - 1342 mB
                     - ZDE crash; to be fixed during the next two contacts
                   AMCS: steady

  15:11:30   WHM5  PSPC: 1465 mB; ZDE: partly ok; AMCS: steady.

  16:51:45   WHM6  AMCS: steady.
                   PSPC: 
                     - 1364 mB (not connected to tanks between WHM 5 and 6)
                     - ZDE: ok
                         'SAFEB' still missing; 
                            ZDE crashed last time when 'SAFEB' was uplinked

  21:34:56   WPS   PSPC:
                     - pressure in A / B / C / PSPC: 1842 / 1046 / 1824 / 1476 mB
                     - All remaining FI uplinks have been sent.


Tue 1998-12-08 (DoY 342)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC is under pressure, and in off position.
    WFC is in safemode.

    'SAFEB' will be replaced by BELT commands (during Bright Earth Blocks)

  08:24:16   WHM1  AMCS: 
                     - safemode!
                     - step 1 of recovery not completed (due to FI commanding)
                   PSPC:
                     - ZDE crashed
                     - pressure in A / B / C / PSPC: 1822 / 1228 / 1824 / 1130 mB
                     - filling of PSPC stopped since last recovery

  10:03:30   WHM2  AMCS: recovery step 1 completed
                   PSPC:
                     - pressure in A / B / C / PSPC: 1832 / 1254 / 1824 / 1099 mB
                     - filling of PSPC started again
                     - ZDE cold start

  11:43:15   WHM3  AMCS: recovery step 2 completed
                   PSPC:
                     - pressure in A / B / C / PSPC: 1842 / 1288 / 1814 / 1446 mB
                     - status as before WHM1 this morning

  13:23:00   WHM4  AMCS: on timeline, and steady
                   PSPC:
                     - pressure in A / B / C / PSPC: 1842 / 1305 / 1824 / 1515 mB
                     - HV was switched on
                     - internal calibration source
                     - filter wheel in open position at the end of the contact

    Now the S/C is observing according to the nominal timeline,
    and also the PSPC is operated nominally.

    Some security mechanisms for the PSPC have been disabled.

  15:02:30   WHM5  AMCS: transient (though it should be steady according to TL)
                   PSPC:
                     - pressure in A / B / C / PSPC: 1822 / 1331 / 1814 / 1487 mB
                         PSPC is now at nominal pressure

  16:43:30   WHM6  AMCS: steady
                   PSPC:
                     - pressure in A / B / C / PSPC: 1832 / 1348 / 1824 / 1487 mB
                     - reasonable signals


Wed 1998-12-09 (DoY 343)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC is operated nominally; WFC is in safemode.

  00:38:33   GDS   AMCS: ok
                   PSPC: pressure in A / B / C / PSPC: 1832 / 1409 / 1824 / 1487 mB

  08:16:30   WHM1  AMCS: transient (though it should be steady)
                   PSPC: pressure in A / B / C / PSPC: 1842 /    ? / 1824 / 1487 mB
                     - HV is off !?
                     - filter wheel voltage is off !?
                     - filter wheel is closed.
                   Actions:
                     - switch on filter wheel voltage, and move to CAL position
                     - switch on PSPC HV: internal calibration: ok

    PSPC is operated nominally again.

  09:54:45   WHM2  AMCS: transient (though it should be steady)
                   PSPC: pressure in A / B / C / PSPC: 1822 / 1520 / 1814 / 1487 mB
                     - Master Veto Rate at 4 cts/s (should be more than 40 cts/s)
                     - something is wrong with the PSPC

    Attitude is off timeline since sod 500 (wrong by 75 deg around sun axis).

  11:34:30   WHM3  AMCS:
                     - recovered to timeline; target coordinates: 
                         R.A.=163.66 deg, Dec=-5.81 deg, Rol=247 deg
                   PSPC: pressure in A / B / C / PSPC: 1842 / 1538 / 1814 / 1486 mB
                     - status as in WHM 1 (HV off); reason unknown
                     - not fixed due to attitude recovery

  13:14:15   WHM4  AMCS: steady
                   PSPC: pressure in A / B / C / PSPC: 1832 / 1538 / 1824 / 1486 mB
                     - HV is off !?
                     - filter wheel is closed.
                   Actions:
                     - switch on PSPC HV: internal calibration (4 min): ok
                     - master veto rate at 230 cts/s (ok) e
                     - move filter wheel to 'open' position
                   ? - gas flow has been changed to nominal value (no "wuerg")

  14:53:45   WHM5  AMCS: steady (from time to time)
                   PSPC: pressure in A / B / C / PSPC: 1832 / 1564 / 1824 / 1485 mB
                     - HV is off !?
                     - filter wheel is closed.
                   Actions:
                     - switch on PSPC HV: internal calibration (4 min): ok
                     - move filter wheel to 'open' position

  21:17:23   WPS   PSPC:
                     - command buffer will be reset
                     - new belt commands will be inserted, which are extended
                         by 3 minutes at the beginning and at the end of
                         the nominal timeline belts.


Thu 1998-12-10 (DoY 344)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC is operated nominally; WFC is in safemode.

  08:07:15   WHM1  AMCS: 
                     - transient (though it should be steady)
                     - large updates during the contact
                     - attitude is obviously again off timeline
                   PSPC: pressure unchanged
                     - HV is off !?
                     - filter wheel is closed.
                   Actions:
                     - switch on PSPC HV

  09:46:00   WHM2  AMCS:
                     - 'quick' recovery performed
                     - smaller updates during the contact
                   PSPC: pressure unchanged
                     - HV is off again, and filter wheel is closed.
                     - switch on PSPC HV; open filter wheel

  11:26:00   WHM3  AMCS: earth block
                   PSPC: pressure unchanged

  13:05:30   WHM4  AMCS: steady at the end
                   PSPC: pressure unchanged
                     - HV ok, filter wheel is open
                     - master veto rate is too low

  14:45:00   WHM5  AMCS: steady
                   PSPC: pressure unchanged
                     - HV ok, filter wheel is open at the beginning
                     - countrates too low after a while
                   ZDE: dump of interrupt adresses

  19:29:32   WPS   AMCS: no safemode

    Suspicion: magnetometer updates are not enabled; check!


Fri 1998-12-11 (DoY 345)

    AMCS is in 'degraded pointing mode'; S/C should be on nominal timeline.
    PSPC is operated nominally; WFC is in safemode.

  07:58:15   WHM1  AMCS: transient (though it should be steady)
                   PSPC: pressure in A / B / C / PSPC: 1832 / 1745 / 1814 / 1484 mB
                     - HV is off for unknown reasons
                     - switch on PSPC HV
                     - bright X-ray source was visible on the quicklook image


    Attitude was off by 50 deg during WHM1.

  09:37:15   WHM2  AMCS: transient
                     - magnetometer updates were enabled
                     - several large (more than 10 deg) updates during the contact
                     - on target at the end of the contact (though not steady)
                   PSPC: pressure in A / B / C / PSPC: 1832 / 1754 / 1814 / 1485 mB
                     - HV is off for unknown reasons; switch on PSPC HV
                     - Master Veto rate ok after switch on
                   GSOC: computers on ground crashed; therefore limitations 
                           in commanding

    Attitude recovery is scheduled for next contact.

  11:17:15   WHM3  AMCS: steady (!); no recovery required
                   PSPC: pressure unchanged
                     - HV is again and again switched off for unknown reasons
                     - internal calibration; seems to be ok

  12:56:45   WHM4  AMCS: steady
                   PSPC: pressure unchanged
                     - HV off at the beginning of the contact
                     - no longer reduction of gas flow
                     - source and hot spot / hot wires visible at the center
                         of the PSPC

    No DSN contacts during the following night (due to shuttle requirements).
    HV will be switched on automatically from time to time.


Sat 1998-12-12 (DoY 346)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC is operated nominally; WFC is in safemode.

    AMCS was steady during the first two contacts.
    Accumulated quicklook image: 
      Hot wire at the center, hot spot at the left of the image.

  11:08:30   WHM3  AMCS: transient and Earth block (as expected)
                   PSPC: pressure in A / B / C / PSPC: 1832 / 1780 / 1814 / 1484 mB
                     - HV is off /filter wheel closed for unknown reasons (PSPC safemode)
                     - switched on PSPC HV, opened filter wheel (PSPC recovery)
                     - bright Earth was observed; 
                         quicklook image and countrates seem to be ok

  12:48:00   WHM4  AMCS: mostly transient
                   PSPC: pressure unchanged
                     - again PSPC safemode; recovery performed
                     - quicklook image looks ok

  14:27:45   WHM5  AMCS: steady
                   PSPC: pressure unchanged
                     - ok (no safemode, and countrates ok)

    Belt extension by 3 minutes don't seem to be necessary;
    PSPC safemodes are rather correlated with slews 
    (bright Earth during slew?).

    Starting on 1998-12-14 noon, the PSPC will be switched on automatically
    towards the end of slews.


Mon 1998-12-14 (DoY 348)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC is operated nominally; WFC is in safemode.

    WFC will be recovered today.

  07:31:30   WHM2  AMCS: steady
                   PSPC: pressure unchanged
                     - HV was ok at the beginning of the contact
                     - countrates dropped to unusual low values during the contact

  09:11:00   WHM3  AMCS: Earth block (as expected)
                   PSPC: pressure unchanged
                     - countrates at unusual low values during the contact

  14:10:30   WHM6  AMCS: Earth block (as expected)
                   PSPC: pressure unchanged
                     - internal calibration
                     - MV and countrates ok


Tue 1998-12-15 (DoY 349)

    AMCS is in 'degraded pointing mode'; S/C should be on nominal timeline.
    PSPC is operated nominally.
    WFC is operated nominally (sun shutter is still closed).


  07:22:30   WHM2  AMCS: transient (though it should be steady)
                     - NRTA: Timeline disrupted since
                   PSPC: pressure unchanged
                     - HV and MV rates are ok
                     - quicklook image: plenty of counts; spectrum ok;
                          strange egg-like structure in the upper half

    An AMCS mini recovery has to be performed as soon as possible.

  09:02:15   WHM3  AMCS: recovery deferred to next contact.
                   PSPC: pressure unchanged
                     - HV and MV rates are ok

    PSPC timeline was not sent for the first part of this day.
    (i.e., PSPC was operated without any protection in radiation belts!)

  10:50:00   WHM4  AMCS: recovery failed
                   PSPC: pressure unchanged
                     - HV and MV rates are ok

  12:21:30   WHM5  AMCS: mini-recovery ok
                   PSPC: pressure unchanged 
                     - HV and MV rates are ok
                     - internal calibration performed

    The first ROSAT night after the full sun orbit period will start at
    13:22:00 (duration according to the timeline: 45 s).

  14:02:00   WHM6  AMCS: back on timeline (Earth block)
                   PSPC: pressure unchanged
                     - HV and MV rates are ok
                     - internal calibration performed

    Due to the mini safemode the WFC has to be recovered again.
    This is scheduled for the next two DSN contacts.

  20:47:36   ----  AMCS: safemode! (sun on the back of ROSAT)

    PSPC Timeline:

      Starting on 1998-12-16, 14:00 UT, the PSPC timeline is generated
      according to the following rules:

        - Belt commands during Belt, Slew, and Earth Block (both bright and dark)
        - Slew durations are shortened by 300 s as long as there are no belts or
            Earth blocks.
        - Belts, Slews, and Earth blocks are merged to one Belt command
            if possible.
        - A PSPC recovery is inserted every 3600 s after a belt command.

Wed 1998-12-16 (DoY 350)

    AMCS is in safemode.
    PSPC is operated nominally.
    WFC is in safemode.

  07:13:45   WHM2  AMCS: complete recovery (within one contact; impressive!)
                   PSPC: pressure unchanged
                     - MV rates are ok; HV off due to Earth block

  08:53:45   WHM3  AMCS: Earth block
                   PSPC: pressure unchanged
                     - HV and MV rates are ok
                     - internal calibration performed

  10:33:30   WHM4  WFC: fully recovered
                     - sun shutter has been opened

  12:12:45   WHM5  AMCS: steady (!)
                   PSPC: pressure unchanged
                     - HV and MV rates ok at the beginning
                     - quicklook image: AR Lac visible
                         at the center of the image;
                         nearly no elongation

  13:54:00   WHM6  AMCS: Earth block
                   PSPC: pressure unchanged
                     - HV off (due to Earth block)

  16:57:38   WPS   AMCS: Earth block in the beginning
                     - update by roughly 2 deg few minutes after Earth block end
                     - steady at the end of the contact
                   PSPC: pressure unchanged
                     - HV and MV ok after end of Earth block


Thu 1998-12-17 (DoY 351)

  00:46:02   ----  AMCS: safemode! (no sun presence in CSS3 for more than 600 s)

    AMCS is in safemode.
    PSPC is operated nominally.
    WFC is in safemode.

  08:45:00   WHM3  AMCS: Earth block
                   PSPC: pressure unchanged
                     - HV and MV are ok; internal calibration

  10:24:45   WHM4  AMCS: Earth block - transient - steady at the end
                   PSPC: pressure unchanged
                     - HV and MV are ok; quicklook image ok

  12:04:00   WHM5  AMCS: Earth block
                   PSPC: pressure unchanged
                     - HV timeout due to Earth block

  18:27:47   WPS




Fri 1998-12-18 (DoY 352)

    AMCS is in 'degraded pointing mode'; S/C is on nominal timeline.
    PSPC and WFC are operated nominally.

  08:36:15   WHM3  Switch off ROSAT, part I
                   S/C: Last tape dump
                   PSPC: pressure in A / B / C / PSPC: 1842 / 1806 / 1835 / 1484 mB
                     - filter wheel closed
                     - carroussel in safe position
                     - ZDE off
                     - gas system:
                         * tank A and C stay connected (valves in between stay open)
                         * valve between tanks and PSPC is closed
                         * valve between PSPC and space is opened
                   WFC: saved
                   AMCS: Earth block, then transient;
                     - finally put into safemode

  10:15:45   WHM4  Switch off ROSAT, part II
                     - receiver on ROSAT stays on
                     - switch tape drives into stand-by
                     - currents as since 980428
  11:16:31           - reset of spacecraft clock
  11:21:22           - transmitter on ROSAT were switched off


========================================================================

         E n d   o f   t h e   R O S A T   M i s s i o n   ( ? )

========================================================================



Remaining tasks:

    Spacecraft:
      - Battery reconditioning
          Concern: battery could fail before the next full 
                   sun orbit period in May 1999

    PSPC:
      - Use rest of gas for a scientific program.

    HRI:
      - Observe strong source with the HRI
      - Door test (strongly requested by the US)

    WFC:
      - Observe AR LAC for at least two days.
      - HZ43 calibration.
      ? Engineering tests.




eof